Using the handshaking principle,determine the number edges of a graph with fourteen vertices and each with degree six


1
Expert's answer
2021-05-07T09:44:26-0400

Find the sum of all degrees of vertices in the graph:

"\sum deg\;v = 6*14=84"

By handshaking lemma,

"\sum deg\;v = 2 \cdot |E|"

Then the number of edges in the graph is

"|E| = \frac{{84}}{2} = 42"

Answer: 42